i need help with row operations the problem is 4y=2x+4
-x=-2-y

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

i need help with row operations the problem is 4y=2x+4
-x=-2-y
---------
Rearrange::
x - y = 2
2x-4y = -4
-------------
Subtract 2 times the 1st row from the 2nd row to get::
x - y = 2
0 -2y = -8
---------
Divide the 2nd row by -2 to get::
x - y = 2
0 + y = 4
----------
Add the 2nd row to the 1st to get:
x + 0 = 6
0 + y = 4
--------
Solution:
x = 6 ; y = 4
===================
Cheers,
Stan H.
====================

Answer by richwmiller(17219)

-x=-2-y
x=2+y
2x-4y=-4
2*(2+y)-4y=-4
4+2y-4y=-4
-2y=-8
y=4
x=2+4
x=6
x = 6, y = 4
x-y=2
2x-4y=-4
multiply by2
2x-2y=4
subtract
-4y+2y=8
2y=8
y=4 etc
1 -1 2
2 -4 -4
2 -2 -4
0 2 8
0 1 4
1 -1 2
1 0 6
0 1 4